OpenStudy (anonymous):

If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)].

OpenStudy (anonymous):

f [g(x)] = 3 ( 2x - 4) + 1 = 6x - 12 + 1 = 6x - 11 -> f[g(2)] = 6*2 - 11 = 1